Go to Planning Wiki - MB Wiki - MB Web - MB Meetings - MB Minutes - MB Action List

2009: Action List for the WLCG MB

2009: Issues to Follow

Collection of usual issues to follow

  • SRM 2.2 deployment
  • HL Milestones
  • SAM tests: NDGF and OSG tests
  • VO Boxes SLA and 24x7 Documents
  • Job Priorities testing, deployment, i/f to batch systems
  • Storage Accounting, Automatic Accounting
  • Tier-2 Reliability and Accounting
  • Pilot Jobs agreement and deployment
  • CPU Benchmarking/HEPiX
  • CPU Efficiency
  • Monitoring and GridView

2009: Actions Done

Closed at MB meeting on 23.3.2010. Names have been provided by those Experiments and Sites willing to contribute.

  • Gstat data: Tier-1 Sites should explain the differences between pledges, monthly accounts and the gstat values (see presentation at the MB 1.12.2009). They should also check their Tier-2 Sites.

The table should be generated every month ad distributed to the Sites. Action removed.

  • I.Bird will prepare the answer to the Scrutiny group about their request to have the 2011 Experiments' Requirements by early 2010.
Done.

  • Each Site should send the URL of the XML File with the Tape Metrics to A.Aimar.

See http://sls.cern.ch/sls/service.php?id=WLCG_Tier1_Tape_Metrics
Sites should send URLs to existing information until they do not provide the required information.

  • 5 May 2009 - CNAF completes, and sends to R.Wartel, their plans to solve the issues concerning security procedures, alarm contact point, etc.

Removed. To send a more complete report.

  • The MB sends to I.Bird input on topics and priorities for the Technical Forum.

  • Sites will be asked to report on the items resulting at the STEP09 Workshop. STEP09_Actions.pdf

  • A.Aimar will contact T.Bell for MSS real-time metrics.

  • Experiments send to A.Sciabá their ranking for the features in the SRM MoU addendum.

  • A.Aimar finds how to display directly SLS information from all Sites, without using the SLS interface, for July’s F2F Meeting. And also which metrics Sites are currently displaying.

Done. Examples from A.Di Girolamo, where the same metric for all Sites is aggregated in a single web page.
http://voatlas08.cern.ch/storage/atlas/utils/storage_space/Storage_Space_ATLASMCDISK.html
https://project-atlas-ddm-sls.web.cern.ch/project-atlas-ddm-sls/ss_vobox_all/

  • M.Schulz should report about the status of the gLExec patch on passing the environment
Done.

  • 30 Jun 2009 - Sites comment on the ALICE dataflow and rates.
Done. No comments from the Sites.

  • VOBoxes SLAs:
    • CMS: Several SLAs still to approve (ASGC, IN2P3, CERN and PIC).
    • ALICE: Still to approve the SLA with NDGF. Comments exchanged with NL-T1.

Alice has approved the NDGF’s SLA.
The action is considered done for CMS and ALICE and the SLAs approved. If there will be issues it will not be responsibility of the Sites.

  • Tier-1 Sites should start publishing the UserDN information.

Done. J.Gordon asked the portal developers in order to have an anonymous list displayed on the web.

  • 9 Jun 2009 - *Sites should report at the MB whether now, after the GDB presentations, the situation of the data rates is clear.

Data rates not presented by ALICE yet. ALICE DONE.

  • 9 Jun 2009 - Experiments should send to J.Gordon the DNs of the people that can read the details of the users in the CESGA Portal.

Done.

  • A group to prepare the plan of the migration regarding the CPU power published by sites through the Information

Removed.

  • Pledges and Requirements need to be updated to HEP-SPEC

Removed.

  • The CESGA production Portal should be verified. J.Gordon will check and send again the information to the MB on which portal to use (prod or pre-prod).

  • A.Aimar schedules a presentation at the F2F in June. Experiments will explain their VO SAM tests and the tests that are reported in their dashboards.

Done.

  • Status of HEPSPEC benchmarking (licensing, running, etc) at each Tier-1 Site.

Done. Verified at thr F2F Meeting in May.

  • M.Schulz will summarize the situation of the User Analysis WG in an email to the WLCG MB.

Done.

  • Experiments will have to explain their VO SAM tests and the tests that are reported in their dashboards.

Done.

  • Web site at CERN should be set up to store the HEP-SPEC benchmark data from the WLCG Sites.

Done. Now available at: https://hepix.caspur.it/benchmarks

  • M.Schulz will come back to the MB with an explanation of the security issues and next steps with gLExec.
Done at the June's GDB.

  • Sites Tier-1 Sites and main Tier-2 Sites buy the license for the benchmark

Removed. The MB will follow only the ier-1. To be done.

  • Countries should comment on the policy document on user information accounting

Removed. This is more a GDB issue.

  • 26 May 2009 - Sites reply which MSS information they can provide in XML already, and whether can be reported by Experiment.

Done.

  • 26 May 2009 - A.Aimar will see how to extract the information about scheduled downtimes from GOCDB and report it to the MB every month.

Done. Scheduled downtimes are registered in GOCDB and A.Aimar will extract the down-time information from GODDB.

  • 26 May 2009 - ASGC should distribute a detailed plan on setting up the Tier-1 systems

Done. ASGC sent the necessary information to I.Bird.

  • I.Bird will look for the chairperson and also distribute a proposal for the mandate of the WLCG Technical Forum and reminding the GDB’s mandate.
Done.

  • 16 Dec 2008 - Sites requested clarification on the data flows and rates from the Experiments. The best is to have information in the form provided by the Data flows from the Experiments. Dataflow from LHCb

Done. The Experiments presented their dataflow and rates at May’s GDB.

    • Convert the current requirements to the new unit.

Done by S.Foffano.

* I.Bird will look for the chairperson and also distribute a proposal for the mandate of the WLCG Technical Forum and reminding the GDB’s mandate.

Done.

  • J.Gordon and R.Pordes will write a requirements document on user information accounting.
Done.

  • 14 Apr 2009 - CNAF reports on how they plan to handle the security incidents report and periodic tests

Done. A new action for L.Dell'Agnello.

  • 4 March 2009 - M.Schulz to present the list of priorities for the Analysis working group.

ACTION REMOVED.
The list of priorities was presented at the WLCG Workshop and then discussed. But the priorities are not fully agreed. R.Pordes volunteered to co-chair the group and will re-open the discussion on the priorities.

  • J.Gordon will send to the GDB the message about the need of 32 compatibility libraries on SL5 64 bits installations.
Not followed by the MB.

  • Sites should report whether (or when) FTS is deployed on SL4.
DONE.

  • A.Aimar will verify with M.Litmaath the situation for the pilot hobs frameworks of ALICE, ATLAS and CMS.

LHCb’s is approved, other have to make modifications. ACTION REMOVED. Already followed monthly at the GDB.

  • 14 Apr 2009 - Sites and Experiments should comment on the need and functions of a WLCG technical group.
DONE.

  • 17 Mar 2009 - Sites should report whether GFAL and lcg-utils can start using by default SRM V2 and will not impact VOs outside WLCG.

DONE.

  • 17 Mar 2009 - Experiments should confirm whether the schedule for the changes regarding “Busy” Storage Services is acceptable.

DONE. Only CMS has replied.

  • SCAS Testing and Certification
M.Schulz reported that after days of stress test there was only one error. All return codes need to be checked and will be released to PPS. Action Completed.

DONE.

  • 17 Feb 2009 - R.Pordes agreed to provide, within 2 weeks, the milestones for OSG reporting installed capacity into APEL.

  • 09 Dec 2008 - Comments to the proposal for collecting Installed Capacity (F.Donno's document) should be sent to the MB mailing list.

Done. Discussed on the 3 February.

  • 3 Feb 2009 - M.Schulz will report about the issues on the SAM servers (and the issues of the ATLAS lock files).

Done. 3 February.

  • 13 Jan 2009 - G. Merino to present a new CPU unit conversion factor proposal to the MB.
DONE.

  • 13 Jan 2009 - Sites present their MSS metrics to the F2F MB.

Removed and will be followed by J.Gordon and discussed at the GDB.

  • DCache teams should report about the client tools should present estimated time lines and issues in providing the porting to gcc 4.3.

Removed. Will be covered by the Architect Forum task on SL5 and gcc 4.3.

  • 6 Jan 2009 - F.Donno, with the help of J.Gordon and R.Pordes, will separate the requirement from the implementations recommended in the document about Reporting Installed Capacity..

Removed. Overlaps with another action.

  • 16 Dec 2008 - P. Mato and M. Schulz to follow-up on the GCC 4.3 discussion, agree on co-ordination and timescale, and report back to the MB.

To remove. Is now followed by the AF and they will keep the MB informed

  • 9 Dec 2008 – I. Bird will check the SRM V2 link in the action list with the SAM team and report back to the MB.

DONE. I.Bird reported that the page reports the AND of all results and the V2 tests.

  • 16 Dec 2008 - Experiments comments the VO-specific Reliability Reports

DONE. Received by email from all Experiments.

  • 09 Dec 2008 - Comments to the proposal of new High Level Milestones should be sent to the MB mailing list.

DONE.

DONE. Same action as the one above about SRM V2.

Done for all Experiments.

  • Experiments should verify whether the gcc 4.1 binaries work and report on issues and problems. P.Mato will report in a couple of weeks.

REMOVED. This action is replaced by the one where M.Schulz and P.Mato agree on a plan for preparation and testing of the SLC5 platform.

  • 30 Nov 2008 - Sites should configure (if needed) their SRM V2 before the end of November. The SRM V2 SAM results are already available in the SAM tests and will be used for the December GridView Reports.

Done.

  • HL Milestones and Metrics
    • The next steps will be to propose some new milestones and target dates.
    • Proposals for metrics are also needed from the Tier-1

Done.

  • F.Donno will distribute a document to describe how the installed accounting is collected and should describe in details the proposed mechanism for sites to publish their inhomogeneous clusters using the current Glue1.3 schema Cluster/Subcluster structures.

Done.

  • A.Aimar should distribute to the MB list where/how the reliability data can be extracted/exported into other formats.

  • 19 Aug 2008 - Feedback from ATLAS on their Job Priorities installations and description of shares required at each Site.
    • For the Job Priorities deployment the following actions should be performed :
      • DONE. A document describing the shares wanted by ATLAS
      • DONE. Selected sites should deploy it and someone should follow it up.
      • ONGOING. Someone from the Operations team must be nominated follow these deployments end-to-end

M.Lamanna and D.Barberis reported ATLAS will prepare a request to WLCG there are 4 user groups that should be recognised by the Sites for production, analysis and local communities. The solution adopted is much simpler than planned originally.

This action can be considered done.

  • P.Mato should report to the MB the progress on the SL5 testing by the Experiments.

Done.

  • Converting Experiments requirements and Sites pledges to new CPU units.

Working group formed.

  • 23 Sept 2008 - Prepare a working group to evaluate possible SAM tests verifying the MoU metrics and requirements.

Is on the agenda for the WLCG Workshop in November.

  • 13 Oct 2008 - O.Keeble will send an updated proposal for software distribution and update at the Tier-1 Sites.

Proposal distributed will be discussed at the GDB.

  • 20 Oct 2008 - By the last week of October the Experiments should provide their new estimated due to the delay in the LHC operations. They should also clarify the assumptions on data taking made in their calculations.

Done. On the 22 October there will be a special meeting to prepare the Overview Board.

  • 8 Oct 2008 - O.Keeble sends to the MB a proposal on possible upgrades to middleware because of the LHC delays.

Proposal distributed by O.Keeble. Approved.

  • Form a working group for User Analysis for a strategy including T1 and T2 Sites.

Done.

  • 9 May 2008 - Milestones and targets should be set for the LCAS solution (deployment on all sites) and for the SCAS solution (development, certification and deployment).

About LCAS: On going. It will be installed on the pre-production test bed PPS at CERN and LHCb will test it. Other sites that want to install it up should confirm it.

About SCAS: The SCAS server seems to be ready and “certifiable” in a week. The client is still incomplete.

DONE.

  • 19 Aug 2008 - New service related milestones should be introduced for VOMS and GridView.
Removed.
  • Experiments QR presentations at the MB. Agreed calendar: ALICE: 30/9, CMS: 30/9, LHCb: 7/10, ATLAS 14/10

  • 18 Set 2008 - Sites should send to A.Aimar and H.Renshall the status of their procurement for the 2009 *installations and whether they are on track for installation by April 2009.

Discussed at the MB of the 16 Sept.

  • M.Schulz should present an updated list of SAM tests for instance testing SRM2 and not SRM1.

Done at September's F2F Meeting.

  • J.Shiers will ask SAM to review the MoU requirements and define SAM tests to verify them.

To remove. A working group will review the MoU looking for which SAM tests are useful (and implementable) to check the MoU requirements.

  • 2 Sept 2008 - B.Panzer will distribute the address of the wiki page with the proposal and comments received about End User Analysis.

Done after the meeting. The wiki page is here: https://twiki.cern.ch/twiki/bin/view/Main/EndUserAnalysisScenario

  • 8 Aug 2008 - O.Keeble should distribute the proposal for the WN distribution (and changes needed at the sites) in writing to the MB and to the GDB.

Done.

  • 31 May 2008 - Sites and Experiments should complete the Operations Alarms Page
https://twiki.cern.ch/twiki/bin/view/LCG/OperationsAlarmsPage

Done.

  • 8 Aug 2008 - WLCG Grid Deployment: all software required at Tier-1 and Tier-2 sites should be described in a wiki page.
Done. https://twiki.cern.ch/twiki/bin/view/LCG/GSSDCCRCBaseVersions

Done.

  • Tier-1 Accounting Report for May to be analysed, corrected and an explanatory email sent to the MB.

Done.

  • 1 July 2008 - The MB should decide on the LHCb's request to use multi-user pilot jobs for testing DIRAC in July.

Accepted.

- MB Older Actions -

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2010-08-05 - DeniseHeagerty
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback